A rectangular auditorium seats 1092 people. The number of seats in each row exceeds the number of rows by 11
Answer by stanbon(75887)

A rectangular auditorium seats 1092 people. The number of seats in each row exceeds the number of rows by 11
------
Let # of rows be "r":
Let # of seats in a row be "s".
-----
Equation:
s = r + 11
sr = 1092
-----
Substitute for "s" and solve for "r"::
r(r+11) = 1092
r^2 + 11r - 1092 = 0
----
Factor::
(r-28)(r+39) = 0
---
r = 28 (# of rows)
s = r+11 = 39 (# of seats in each row)
--------------------
Cheers,
Stan H.
==================